OpenStudy (anonymous):

how much power is needed to lift 200-n object to a height of 4m in 4 s ?

OpenStudy (anonymous):

velocity is 1m/s. W = KEf - KEi. (where f=final and i=initial). KE =1/2mv^2. KEi=0 because initial velocity is 0, so W=KEf=1/2mv^2=1/2*m*1. Just plug in mass

OpenStudy (anonymous):

is 200-n the mass, or a 200 newton object?

OpenStudy (anonymous):

power=work/time work=mgh =200*10*4=8000J power=8000/4=2000W

OpenStudy (anonymous):

newton object

OpenStudy (anonymous):

2000w

OpenStudy (anonymous):

so you got it.